This exercise is based on the following table, which lists interest rates on long-term investments (based on 10-year government bonds) in several countries in 2014.†
Country U.S. Japan Germany. Australia Brazil
Yield 2.5%. 0.5% 1.0% 3.5% 4.2%
Assuming that you invest $13,000 in Japan, how long (to the nearest year) must you wait before your investment is worth $20,000 if the interest is compounded annually?
_______ yr

Answer by Boreal(15235)

20000=13000(1.005)^n, n is number of years
1.538=1.005^n
logs both sides
0.18708=nlog(1.005)
n=86.37 years or 86 years